# Constraint-Aware Feature Learning for Parametric Point Cloud

### v5

@@ Abstract
Parametric point clouds are sampled from CAD shapes and are becoming increasingly common in industrial manufacturing. Most existing CAD-specific deep learning methods only focus on geometric features, while overlooking constraints which are inherent and important in CAD shapes. This limits their ability to discern CAD shapes with similar appearance but different constraints. To tackle this challenge, we first analyze the constraint importance via a simple validation experiment. Then, we introduce a deep learning-friendly constraints representation with three vectorized components, and design a constraint-aware feature learning network (CstNet), which includes two stages. Stage 1 extracts constraint feature from B-Rep data or point cloud based on shape local information. It enables better generalization ability to unseen dataset after model pre-training. Stage 2 employs attention layers to adaptively adjust the weights of three constraints’ components. It facilitates the effective utilization of constraints. In addition, we built the first multi-modal parametric-purpose dataset, i.e. Param20K, comprising about 20K shape instances of 75 classes. On this dataset, we performed the classification and rotation robustness experiments, and CstNet achieved 3.52% and 26.17% absolute improvements in instance accuracy over the state-of-the-art methods, respectively. To the best of our knowledge, CstNet is the first constraint-aware deep learning method tailored for parametric point cloud analysis in CAD domain. Our project page with source code is available at: https://cstnetwork.github.io/.
@@ 1 Introduction
Parametric point clouds are sampled from CAD shape instances or captured from various industrial scenarios by common digital devices. As illustrated in Fig. 1, CAD shape instances are typically instantiated from parametric templates, which comprise primitives, constraints, and shape parameters. By specifying a series of parameter configurations, a template can be instantiated into a family of CAD shape instances [62]. In industrial domain, many CAD shapes are visual similar but have different constraints in their functional regions. This poses challenges for existing point cloud deep learning methods which based on geometric features only. Thus, a constraint-aware feature learning method for parametric point cloud is fundamental for a wide range of industrial applications, such as engineering model design and product manufacturing.
@@
First, most existing studies [58, 41, 26, 65] focus on primitive regularity (e.g. uniform appearance and distinct boundaries). They usually leverage geometric features only, but ignore those constraints between primitives. Although these methods have been effectively applied in point cloud segmentation and primitive reconstruction, they inherently limit their discrimination capability for CAD shapes that are visually similar yet functionally distinct. While some studies discussed constraints in 2D engineering sketches [8, 40], the considered primitives in 2D sketches are supposed available and generally few in number. Thus, these approaches can be hardly extended to 3D parametric point clouds. Second, most CAD-specific methods rely on specialized datasets with additional labels (e.g. primitive, parameter, and normal labels). However, point clouds are typically acquired via LiDAR or RGB-D cameras, and in most cases only contain coordinate information. It thus restricts the practical applicability of these approaches. Third, there are still numerous challenges in constraint-aware point cloud analysis. Constraints represent the spatial relationships between primitives, but the number and types of primitives cannot be directly obtained from point clouds. Predicting primitives from point clouds also poses challenges. The unordered nature of point clouds results in unordered outcomes, leading to a mismatch between the predicted primitives and those in the constraint labels. In addition, processing relations between every pair of primitives would produce an overwhelming amount of data, since some CAD shapes contain many primitives. Assuming that the constraints have been obtained, effectively leveraging them remains a significant challenge.
@@
To address the aforementioned issues, we transform the traditional constraint representation (e.g. constraints in SolidWorks or UG [3]) into three vectorized sub-features: primitive’s Main Axis Direction (MAD), Adjacency (Adj), and Primitive Type (PT). Given the completeness of traditional constraint representation, these three sub-features can effectively capture constraints. Representing constraints with MAD, Adj, PT converts the relations between primitive pairs into relations between primitives and a common reference, thereby significantly reducing the total number of relations. Moreover, the MAD, Adj, PT can be inferred by local point patches, and the local features of CAD shapes are predominantly composed of primitives such as planes and cylinders. Consequently, a model that relies solely on local features can generalize to unseen datasets after pre-trained, this property eliminates the reliance on constraint labels during practical applications. To mitigate the impact of point cloud unordered nature, we encode MAD, Adj, PT as point-wise features. As a result, each point with constraints in point cloud is represented as (X, Y, Z, MAD, Adj, PT). The aforementioned processing transform the constraints into deep learning-friendly form. After that, we designed the constraint prediction from point cloud (CstPnt) module to obtain MAD, Adj, PT. The CstPnt module is based solely on local features, experiments show that it exhibits strong generalization ability on unseen datasets, and its training data can be easily obtained by processing unlabeled B-Rep data. Following constraint acquisition, we separately extract features from MAD, Adj, PT, and apply attention mechanism to weight these features, thereby facilitating their effective utilization.
@@
We quantitatively evaluate CstNet on ABC [20] and Param20K dataset: constraints prediction, classification, and rotation robustness evaluation. For constraints prediction, our CstPnt module (in CstNet Stage 1) surpasses the state-of-the-art (SOTA) by around 5% and 14% instance accuracy in ABC (test set) and Param20K (unseen dataset), respectively. For classification and rotation robustness on Param20K, CstNet outperforms the SOTA by 3.52% and 26.17% instance accuracy, respectively.
@@
  • We introduced a deep learning-friendly vectorized representation and designed the first constraint-aware feature learning network CstNet for effective constraint acquisition and leveraging for parametric point cloud.
  • We built the first parametric-purpose Param20K dataset, which consists of about 20K shape instances of 75 classes. Each instance include mesh, point cloud, and B-Rep data. Given the current scarcity of labeled CAD shape datasets, Param20K offers a solid basis for future studies in parametric CAD domains.
  • We conducted extensive experiments on constraint acquisition, classification, robustness evaluation and ablation studies. The performance of CstNet has a significant improvement compared with SOTA methods which usually focus on geometric data only.
@@ 2 Related Work
General point cloud analysis. View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [42, 17, 7, 60, 59, 47]. How to address information loss during the projection and optimizing viewpoints play a critical role. Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [33, 52, 39, 45, 19, 22]. The size and orientation of 3D voxels significantly influences the accuracy of these methods, and information loss occurs during the conversion to voxels. Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [36, 37, 27, 12, 57, 2, 43, 30, 64, 66, 11, 9, 55, 54, 53, 38, 31, 29, 46, 51, 5, 63]. The unordered nature and rotational invariance of point clouds make it challenging to employ methods similar to traditional CNNs for feature extraction, and kNN is a commonly adopted approach.
@@
Other CAD-specific studies. B-Rep-based methods typically involve B-Rep data conversion and feature extraction, the challenge lies in converting B-Rep into deep learning compatible format. These approaches are commonly applied to classification [15], segmentation [21, 6], assembly [16, 48], engineering sketch generation [13], and CAD operation sequence prediction [50, 67]. Mesh-based learning methods use triangle models as input, such as STL and OBJ, contributions often rely on mechanical datasets provided by the authors. These methods are primarily used for classification [32], retrieval [4], segmentation [44], and B-Rep generation [10]. Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [10, 49, 34, 40, 28], 3D models [23, 24, 62], or CAD operation sequences [56]. A key challenge is incorporating the regularity and connectivity of primitives, which is typically overlooked in graphic object analysis.
@@ 3.1 Is Constraint Important?
We aim to utilize constraints to distinguish visually similar but functionally distinct CAD shapes. To answer whether constraints could enhance deep learning methods’ performance, we designed the following validation experiments. The objective is to compare the prisms shown in Fig. 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them. The constraint-aware model is built on PointNet++ backbone, which predicts constraints and then adopts it for classification. The constraints are represented as the relation between the point attached primitive and a reference plane, more details in supplementary.
@@
However, the constraint learning method designed in the validation experiments is only applicable to simple scenarios. To develop a more generalized approach, we further investigated the constraint representation and deep learning architecture tailored for parametric point clouds.
@@ 3.2 Constraint Representation
In general, constraints should be defined as point-wise attribute for point clouds, analogous to normal vector. As point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation. Following the constraint definitions in CAD software is not applicable. However, once the constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraints on CAD shapes can be defined comprehensively.
@@
The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding. The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Fig. 4, relations such as parallel and vertical could be derived from the direction of their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Thus, additional data is required to supplement this information. Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives. For example, in a cuboid, the adjacency and normal of the six faces ensure that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency is combined with main axis directions, the positional relation can be determined comprehensively. Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector. In summary, each composition of constraints in parametric point cloud is illustrated in Fig. 5, and each point with constraint is expressed as (X, Y, Z, MAD, Adj, PT).
@@
Figure 4: Main Axis Direction of primitives. Plane: normal vector. Cylinder, cone, sphere: rotation axis. Main Axis Direction requires unique process, details are provided in Supplementary.
@@
We represent constraints by MAD, Adj, PT, is this approach sufficient to handle most constraints? For PT, commercial CAD systems typically classify primitives into plan, cylinder, cone, sphere, and free-form surface. Among these, plane, cylinder, and cone are primarily influenced by constraints, while spherical and free-form surfaces rely heavily on their geometric features. Therefore, considering plane, cylinder, and cone is sufficient for constraint representation. For sphere and free-form surface, the Stage 2 of CstNet also incorporate geometric features (feed in coordinates). MAD correspond to directional constraints (e.g. parallel and vertical), and MAD combined with PT represent dimensional constraints (e.g. distance and tangent). Through MAD and Adj, most common constraint types can be covered. Therefore, the combination of MAD, Adj, and PT is sufficient to represent the majority of constraints.
@@ 3.3 Constraint Acquisition
The Stage 1 of CstNet is designed to obtain constraints, as shown in the left of Fig. 7. When B-Rep data is available, the constraint computation from B-Rep (CstBRep) module is developed to obtain constraints. For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [25] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped. For adjacency computation, we identify valid edges firstly. An edge is considered invalid if the two adjacent primitives are smoothly connected along the edge. Searching valid edges is accomplished by OCCT. Next, we compute the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (set to $0.08\times ShapeArea/4\pi$ in this paper). For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
@@
In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted to point cloud, after which the CstPnt module can be used to predict constraints. Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary. The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps: Step 1: Identify neighbor points around $\boldsymbol{p}$. Step 2: From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, store in $\boldsymbol{\mathrm{P_{valid}}}$. Step 3: Fit shapes such as cylinders or planes using $\boldsymbol{\mathrm{P_{valid}}}$, and determine the primitive type with the smallest fitting error. Step 4: Based on $\boldsymbol{\mathrm{P_{valid}}}$ and the primitive type, calculate the main axis direction and assess whether the point is near an edge. While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
@@
The design of CstPnt module is inspired by the above steps, with details in Fig. 7. For Step 1, the traditional approach is to use the k-Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere. However, for constraint computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig. 6. To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary. For Step 2, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding, the relation between points is designed as embedding subtraction, as shown in Eq. 1. Step 3 and Step 4 involve computing MAD and PT, both steps can be accomplished by MLPs. We employed a U-Net-like structure for better data utilization.
@@
Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication. $\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in similar structure by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$, and then feeding it into MLPs, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
@@
The structure of CstPnt reveals that only local information is utilized, which aligns with the local nature of the constraint representation we designed. For no global information is incorporated and most CAD shapes consist of a limited set of primitives, such as plane, cylinder, and sphere, the CstPnt could generalize to unseen datasets if pre-trained on a large scale dataset. The training data is derived from unlabeled B-Rep data, as a result, our CstPnt enables the effective use of existing large scale unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
Figure 7: Overview of CstNet. The top side illustrates the overall architecture of CstNet, which comprising two stages. Stage 1 is designed for constraint acquisition. When B-Rep data is available, the CstBRep module is utilized to extract constraints; otherwise, use the CstPnt module. Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes. The bottom side presents the details of module design. In the CstPnt module, the final MLP consists of three parallel MLPs with the same input, where the outputs correspond to Main Axis Direction, Adjacency, and Primitive Type.

### v6

@@ Abstract
Parametric point clouds are sampled from CAD shapes and are becoming increasingly common in industrial manufacturing. Most CAD-specific deep learning methods focus on geometric features, while overlooking constraints inherent in CAD shapes. This limits their ability to discern CAD shapes with similar appearances but different constraints. To tackle this challenge, we first analyze the constraint importance via simple validation experiments. Then, we introduce a deep learning-friendly constraints representation with three components, and design a constraint-aware feature learning network (CstNet), which includes two stages. Stage 1 extracts constraint representation from BRep data or point cloud based on local features. It enables better generalization ability to unseen dataset after pre-training. Stage 2 employs attention layers to adaptively adjust the weights of three constraints’ components. It facilitates the effective utilization of constraints. In addition, we built the first multi-modal parametric-purpose dataset, i.e. Param20K, comprising about 20K CAD instances of 75 classes. On this dataset, CstNet achieved 3.49% (classification) and 26.17% (rotation robustness) accuracy improvements over the state-of-the-art. To the best of our knowledge, CstNet is the first constraint-aware deep learning method tailored for parametric point cloud analysis. Our project page with source code is available at: https://cstnetwork.github.io/.
[…]
@@ 1 Introduction
Parametric point clouds are sampled from CAD shapes or captured from various industrial scenarios by common digital devices. As illustrated in Fig. 1, CAD shapes are typically instantiated from parametric templates, which comprise primitives, constraints, and shape parameters. By specifying a series of parameter configurations, a template can be instantiated into a family of CAD shapes [67]. In industrial domain, many CAD shapes are visual similar but have different constraints in their functional regions, examples shown in Fig. 1. This poses challenges for geometric features-based point cloud deep learning methods. Thus, a constraint-aware feature learning method for parametric point cloud is fundamental for a wide range of industrial applications, such as engineering model design [67, 18].
@@
First, most existing studies [63, 43, 28, 70] focus on primitive regularity (e.g. uniform appearance and distinct boundaries). They usually leverage geometric features only, but ignore constraints between primitives. Although these methods have been effectively applied in point cloud segmentation and primitive reconstruction, they inherently limit their discrimination capability for CAD shapes that are visually similar yet functionally distinct. While some studies discussed constraints in 2D engineering sketches [10, 42], the considered primitives in 2D sketches are supposed available and generally few in number. Thus, these approaches can be hardly extended to 3D parametric point clouds. Second, most CAD-specific methods rely on specialized datasets with additional labels (e.g. primitive, parameter, and normal labels). However, point clouds are typically acquired via LiDAR or RGB-D cameras, and in most cases only contain coordinate information. It thus restricts the practical applicability of these approaches. Third, there are still numerous challenges in constraint-aware point cloud analysis. Constraints represent the spatial relationships between primitives, but the number and types of primitives cannot be directly obtained from point clouds. Predicting primitives from point clouds also poses challenges, the point cloud unordered nature results in unordered outcomes, leading to a mismatch between the predictions and labels. In addition, processing relations between every primitive pairs would produce an overwhelming data volume, since some CAD shapes contain many primitives. Assuming that the constraints have been obtained, effectively leveraging them remains a significant challenge.
@@
To address the aforementioned issues, we transform the traditional constraint representation (e.g. constraints in SolidWorks or UG [4]) into three components: primitive’s Main Axis Direction (MAD), Adjacency (Adj), and Primitive Type (PT). Representing constraints with MAD-Adj-PT converts the relations between primitive pairs into relations between primitives and a common reference, thereby significantly reducing the data volume. Moreover, MAD-Adj-PT can be inferred by local point patches, and the local features of CAD shapes are predominantly composed of primitives such as planes and cylinders. Consequently, a model that relies solely on local features can generalize to unseen datasets after pre-trained, this property eliminates the reliance on constraint labels during practical applications. To mitigate the impact of point cloud unordered nature, we encode MAD-Adj-PT as point-wise features. As a result, each point with constraint representation in point cloud is defined as $(x,y,z,MAD,Adj,PT)$. The aforementioned processing transform the traditional constraint representation into deep learning-friendly form. After that, we designed the Constraint-Aware Feature Learning Network (CstNet), which consist of two stages. Stage 1 is designed to obtain MAD-Adj-PT. The architecture design is based solely on point cloud local features, experiments show that it exhibits strong generalization ability on unseen datasets, and its training data can be easily obtained by processing unlabeled BRep data. Following MAD-Adj-PT acquisition, Stage 2 separately extract features from MAD-Adj-PT, and apply attention layers to weight these features, thereby facilitating their effective utilization.
@@
We quantitatively evaluate CstNet on ABC [22] and Param20K dataset: MAD-Adj-PT prediction, classification, and rotation robustness evaluation. For MAD-Adj-PT prediction, CstNet surpasses the state-of-the-art (SOTA) by around 5% and 14% instance accuracy in ABC (test set) and Param20K (unseen dataset), respectively. For classification and rotation robustness experiments on Param20K, CstNet outperforms the SOTA by 3.49% and 26.17% instance accuracy, respectively.
@@
  • We introduced a deep learning-friendly constraint representation and designed the first constraint-aware feature learning network CstNet for effective constraint acquisition and leveraging for parametric point cloud.
  • We built the first parametric-purpose Param20K dataset, which consists of about 20K CAD instances of 75 classes. Each instance include mesh, point cloud, and BRep data. Given the current scarcity of labeled CAD shape datasets [20, 16], Param20K offers a solid basis for future studies in parametric CAD domains.
  • We conducted extensive experiments on constraint acquisition, classification, robustness evaluation and ablation studies. The performance of CstNet has a significant improvement compared with SOTA which usually focus on geometric features only.
@@ 2 Related Work
General point cloud analysis. View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [44, 19, 9, 65, 64, 51]. How to address information loss during the projection and optimizing viewpoints plays a critical role. Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [36, 57, 41, 49, 21, 24]. The size and orientation of 3D voxels significantly influences the accuracy of these methods, and information loss occurs during the conversion to voxels. Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [38, 39, 29, 14, 62, 3, 47, 33, 69, 71, 13, 11, 60, 59, 58, 40, 34, 32, 50, 55, 6, 68, 7]. The unordered nature and rotational invariance of point clouds make it challenging to employ methods similar to traditional CNNs for feature extraction, and kNN is a commonly adopted approach.
@@
Other CAD-specific studies. BRep-based methods typically involve BRep data conversion and feature extraction, the challenge lies in converting BRep into deep learning compatible format. These methods are commonly applied to classification [17], segmentation [23, 8], assembly [18, 52, 66], engineering sketch generation [15], and CAD operation sequence prediction [54, 72]. Mesh-based methods use triangle models as input, such as STL and OBJ, contributions often rely on mechanical datasets provided by the authors. These methods are primarily used for classification [35], retrieval [5], segmentation [48], and BRep generation [12]. Sketch-based learning takes free-hand sketches as input, with outputs typically being engineering sketches [12, 53, 37, 42, 31], 3D models [25, 26, 67], or CAD operation sequences [61, 45, 46]. A key challenge is incorporating the regularity and connectivity of primitives, which is typically overlooked in graphic object analysis.
@@ 3.1 Is Constraint Important?
To answer whether constraints could enhance deep learning methods’ performance, we designed the following validation experiments. The objective is to discriminate prisms shown in Fig. 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it more challenging to distinguish them. The constraint-aware model is built on PointNet++ [39] backbone, which predicts constraints and then adopts it for classification. The constraints are represented as the relation between the point attached primitive and a reference plane, more details in $Suppl.$
@@
However, the constraint-aware learning method designed in the validation experiments is only applicable to simple scenarios. To develop a more generalized approach, we further investigated the constraint representation and deep learning architecture tailored for parametric point clouds.
@@ 3.2 Constraint Representation
In general, constraint representation should be defined as point-wise attribute for point clouds, analogous to normal vector. As point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation. Following the constraint definitions in CAD software is not applicable. However, once the constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraint representation on CAD shapes can be defined.
@@
The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding. The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Fig. 4, relations such as parallel and vertical could be derived from their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives. For example, in a cuboid, the adjacency and normal of the six faces ensure that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency is combined with main axis directions, the positional relation can be determined comprehensively. Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by one-hot vector. In summary, constraint representation in parametric point cloud is illustrated in Fig. 5, and each point with constraint is expressed as $(x,y,z,MAD,Adj,PT)$.
@@
Figure 4: MAD of primitives. Plane: normal vector. Cylinder, cone, sphere: rotation axis. MAD computation requires unique process, details are provided in $Suppl.$
@@
Is MAD-Adj-PT sufficient to handle most constraints? For PT, commercial CAD systems typically classify primitives into plane, cylinder, cone, sphere, and free-form surface. Among these, plane, cylinder, and cone are the most common primitives, e.g., for a Trunk of ABC [22], plane: 69.63%, cylinder: 17.51%, cone: 6.99%, other : 5.87%. Besides that, constraints usually applied to regular primitives, while free-form surfaces rely heavily on geometric features. Therefore, considering plane, cylinder, and cone is sufficient for common constraint representation. For free-form surfaces, the CstNet Stage 2 also incorporates geometric features. MAD corresponds to directional constraints (e.g. parallel, vertical), and MAD combined with PT can represent dimensional constraints (e.g. distance, tangent). Through MAD and Adj, most common constraint types can be covered. Therefore, the combination of MAD, Adj, and PT is sufficient to represent the majority of constraints.
@@ 3.3 Constraint Acquisition
The Stage 1 of CstNet is designed to obtain constraint representation, as shown in the left of Fig. 7. When BRep data is available, the constraint computation from BRep (CstBRep) module is selected. For the input BRep data, OCCT [2] performs tessellation to generate a mesh model, followed by Poisson disk sampling [27] via pymeshlab to ensure a uniform point distribution and control the point count. If point cloud generation is not needed, this step could be skipped. For Adj, OCCT identify valid edges firstly. An edge is considered invalid if the two adjacent primitives are smoothly connected along the edge. Next, computing the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (we use $0.08\times ShapeArea/4\pi$). For the PT and MAD, OCCT first determine which primitive it attached, and then analyze the primitive, providing both the PT and MAD.
@@
In the absence of BRep data, such as working with mesh or voxel models, it can be converted to point cloud, after that the constraint prediction from point cloud (CstPnt) module can be used. Since each point’s MAD-Adj-PT are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary. The MAD-Adj-PT computation of a point $\boldsymbol{p}$ could be divided into the following steps: Step 1: Identify neighbor points around $\boldsymbol{p}$. Step 2: From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, store in $\boldsymbol{\mathrm{P_{valid}}}$. Step 3: Fit shapes such as cylinders or planes using $\boldsymbol{\mathrm{P_{valid}}}$, and determine the primitive type with the smallest fitting error. Step 4: Based on $\boldsymbol{\mathrm{P_{valid}}}$ and the primitive type, calculate the main axis direction and assess whether the point is near an edge. While the above process can be carried out by traditional methods, it is highly complex, therefore CstPnt is designed to achieve the same.
@@
The design of CstPnt module is inspired by the above steps, with details in Fig. 7. For Step 1, the k-Nearest Neighbors (KNN) algorithm is commonly used, which identifies neighbors within a sphere. However, for MAD-Adj-PT computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig. 6. To this end, we proposed a neighbor searching method SurfaceKNN, details in $Suppl.$ For Step 2, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding, as shown in Eq. 1. Step 3 and Step 4 can be accomplished by MLPs, we employed a U-Net-like structure for better data utilization.
@@
Where $\boldsymbol{f}_{\mathit{i}}$ is the feature of $\boldsymbol{p}_{i}$, $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication. $\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in similar structure by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$, and then feeding it into MLPs, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
@@
The structure of CstPnt reveals that only local information is used, which aligns with the local nature of the constraint representation we designed. Since no global information is incorporated and most CAD shapes consist of a limited set of primitives, such as plane, cylinder, and cone, the CstPnt could generalize to unseen datasets after pre-training on a large dataset. The training data is derived from unlabeled BRep data, as a result, our CstPnt enables the effective use of existing large unlabeled BRep datasets, making it possible to serve meaningful tasks.
Figure 7: Overview of CstNet. The top side illustrates the overall architecture of CstNet, which comprising two stages. Stage 1 is designed for constraint acquisition. When BRep data is available, the CstBRep module is utilized to extract constraints; otherwise, use the CstPnt module. Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes. The bottom side presents the details of module design. In the CstPnt module, the final MLP consists of three parallel MLPs with the same input, where the outputs correspond to Main Axis Direction, Adjacency, and Primitive Type.